OpenStudy (anonymous):

**will give medal** Write an equation in y = mx + b form for the table. Gallons of gas used 2 4 6 8 Miles driven 52 104 156 208 A. y = 26x B. y = 26x + 26 C. y = 26x + 1 D. y = 26x -1

OpenStudy (ybarrap):

Take two points, say <2,52> and <4,104> and find the slope. This will be your m. Next, take any point, say <2,52> again and plug into y=mx +b. this would be 52=2m+b. You know m, now solve for b. You will them have m and b for y=mx+b. That's it, you're done.

undeadknight26 (undeadknight26):

I think this is y = mx + b where B is days and m is miles. aka B.